# thanks: Corresponding author

### v2

@@ Abstract
The charmonium system, bound states of charm and anti-charm quarks, serve as a critical laboratory for exploring non-perturbative quantum chromodynamics (QCD) [1]. It is relatively simple compared to light hadrons – owing to the heavy quark mass, making it an ideal testbed for theoretical frameworks that bridge perturbative and non-perturbative physics.
@@ I Light-front wave functions
where, $p$ is the 4-momentum of the particle; $j$ and $\sigma$ are the particle’s total angular momentum and magnetic projection, respectively. The coefficients $\psi_{s\bar{s}/h}^{\sigma}(x,\vec{k}_{\perp})$ are the valance sector LFWFs, with $s(\bar{s})$ representing the spin of the quark (antiquark). $x=p^{+}_{1}/p^{+}$ is the longitudinal momentum fraction, $\vec{k}_{\perp}=\vec{p}_{1\perp}-x\vec{p}_{\perp}$ is the relative transverse momentum and $p_{1}^{2}=p_{2}^{2}=m_{q}^{2}$. The ellipsis represents the non-valence Fock contributions. The quark and antiquark creation operators satisfy the anticommutation relations:
@@
$\psi_{\uparrow\downarrow-\downarrow\uparrow/P}(x,\vec{k}_{\perp})=\sqrt{\frac{x(1-x)}{2}}\int{\mathrm{d}}^{3}ze^{{\frac{i}{2}}xp^{+}z^{-}-i\vec{k}_{\perp}\cdot\vec{z}_{\perp}}\\ \langle 0|\frac{1}{\sqrt{N_{c}}}\bar{\psi}(0)\gamma^{+}\gamma_{5}\psi(z)|P(p)\rangle_{z^{+}=0,\vec{p}_{\perp}=0},$ (4)
@@
The second term appears because the LFWFs in general depend on the orientation of light front $\omega$, which is a null vector ($\omega^{2}=0$, and $\omega\cdot\gamma=\gamma^{+}$). In some phenomenological applications, the second term is dropped [64, 65, 66, 67, 68, 69, 70]. However, we have shown that this term plays an important role in chiral symmetry breaking [71]. In both BLFQ and DSE, the maximal kinematical symmetries of the LFWFs are retained and both approaches produce two independent structures. Fig. 1 compares the valence LFWFs of $\eta_{c}$ as obtained from BLFQ and from DSE. The LFWFs of BLFQ are obtained in Ref. [21] and accessible from data repository [87]. For both spin-singlet $\psi_{\uparrow\downarrow-\downarrow\uparrow}$ and spin-triplet $\psi_{\uparrow\uparrow}/\psi_{\downarrow\downarrow}$ components, the LFWFs from BLFQ are broader than DSE in the transverse momentum ($k_{\perp}$) direction and longitudinal ($x$) direction. However, the normalizations of each of the spin configurations are qualitatively similar in these two frameworks. The normalization for each spin configurations is defined as
@@
A comment is in order here. The valance Fock sector LFWFs projected from the DSE are not automatically normalized to 1 due to contributions from high Fock states in the BSA. However, for heavy quarkonia studied in this paper, the valance Fock sector contributes more than 90% of the total normalization [54]. To be consistent with the BLFQ approach, we have normalized the DSE LFWFs to 1. All subsequent calculations will use these normalized LFWFs.
@@ II Form factors and hadronic densities
$\langle O\rangle=\sum_{n}\frac{1}{S_{n}}\prod_{i}^{n}\sum_{s_{i}}\int\frac{{\mathrm{d}}x_{i}}{4\pi x_{i}}\int{\mathrm{d}}^{2}r_{i\perp}4\pi\delta(\sum_{j}x_{j}-1)\\ \times\delta^{2}(\sum_{j}x_{j}\vec{r}_{j\perp})\big|\psi_{n}(\{x_{i},\vec{r}_{i\perp},s_{i}\})\big|^{2}O(\{x_{i},\vec{r}_{i\perp},s_{i}\}).$ (10)
Here, $S_{n}$ is the symmetry factor.
For charmonium, the physical charge form factor vanishes due to the charge conjugation symmetry. Nevertheless, it is custom to define a fictitious charge form factor, where the photon couples to the quark and antiquark differently, similar to the charged pion $\pi^{\pm}$ and $B_{c}$, while the quark masses remain unchanged $m_{q}=m_{\bar{q}}=m_{c}$ [78]. Fig. 2 compares the charge form factor computed using LFWFs from BLFQ and from DSE. We adopt $N_{\mathrm{max}}=8$ LFWFs for central values of BLFQ results and quote the difference between $N_{\mathrm{max}}=8$ and $N_{\mathrm{max}}=16$ as the basis sensitivity [21]. The primary source of uncertainty in the DSE results stems from the RL approximation, which is difficult to estimate. Consequently, only the central values are presented here and in the subsequent calculations. As shown in Fig. 2, the BLFQ and DSE results are in good agreement with each other including at high $Q^{2}$ ($Q^{2}\gtrsim 10\,\mathrm{GeV^{2}}$).
@@
$\langle\psi(p^{\prime})|T^{\mu\nu}(0)|\psi(p)\rangle=2P^{\mu}P^{\nu}A(Q^{2})\\ +\frac{1}{2}(q^{\mu}q^{\nu}-g^{\mu\nu}q^{2})D(Q^{2})$ (11)
where, $P=(p^{\prime}+p)/2$, $q=p^{\prime}-p$ and $Q^{2}=-q^{2}$. In light-front dynamics, $A$ and $D$ are related to the Fourier transform of the transverse flow density $\mathcal{A}(r_{\perp})$, and transverse shear density $\mathcal{T}^{12}(r_{\perp})=\mathcal{S}(r_{\perp})/(2P^{+})$
@@
Fig. 3 compares the gravitational form factors $A(Q^{2})$ and $D(Q^{2})$ computed using LFWFs from BLFQ and from DSE. As one can see, the BLFQ and DSE results are in good agreement. The $D$-term, defined as $D\equiv D(0)$, is known to be sensitive to the interaction. In Ref. [83], the authors argue that the $D$-term of the pseudoscalar meson should lie within (1,1/3)(-1,-1/3), which are the values in the chiral limit and the infinite mass limit respectively. Using DSE with a contact interaction, they obtain D=0.58D=-0.58 for $\eta_{c}$. In contrast, the values extracted from BLFQ and from DSE LFWFs in this work are $D_{\textsc{blfq}}=-4.6$ and $D_{\textsc{dse}}=-4.1$.
@@ III Light-cone distribution amplitudes and radiative transitions
$\langle 0|\bar{\psi}(+{\frac{1}{2}}z)\gamma^{+}\gamma_{5}\psi(-{\frac{1}{2}}z)|P(p)\rangle_{z^{+},\vec{z}_{\perp}=0,\mu}\\ =if_{P}p^{+}\int_{0}^{1}{\mathrm{d}}xe^{{\frac{i}{2}}(x-\frac{1}{2})p^{+}z^{-}}\phi_{P}(x,\mu).$ (16)
@@
Fig. 4 compares the leading-twist LCDA of $\eta_{c}$ as obtained from BLFQ and from DSE and evolved to $\mu=2.8\,\mathrm{GeV}$. The DSE result appears narrower than the BLFQ result. However, both are broad enough to distinguish from the nonrelativistic limit. The decay constant can be derived after integrating over $x$ in Eq. (17),
@@
Figure 4: (Colors online) Comparison of the leading-twist LCDA of $\eta_{c}$ as obtained from BLFQ ($N_{\text{max}}=8$, solid lines) and from DSE (dashed lines). The uncertainty bands are computed from the difference between $N_{\text{max}}=8$ and $N_{\text{max}}=16$ results in BLFQ. The dot-dashed lines correspond to the pQCD asymptotic form 6x(1x)6x(1-x) [7]. The DSE results are evolved from $\mu_{0}=2.6\,\mathrm{GeV}$ to $\mu=2.8\,\mathrm{GeV}$ to match the BLFQ energy scale.
The LCDA is an essential tool for describing exclusive processes in high-energy scattering $Q^{2}\ll\Lambda_{\textsc{qcd}}^{2}$. One of the observables that can be described by LCDA is the two-photon transition form factor (TFF), as shown by Lepage & Brodsky [7]. On the other hand, the evaluation of the TFF can be extended to low $Q^{2}$ via LFWFs [7, 85, 27]. Fig. 5 compares the evaluation of the TFF $F_{\eta_{c}\gamma}(Q^{2})$ for the process $\gamma^{*}\gamma\to\eta_{c}$ using LFWFs obtained from BLFQ and from DSE. The theoretical results, without tuning of parameters, are in good agreement with BaBar measurements.
@@ IV Summary and conclusions
This work also validates the reliability of the BSA projection method in heavy quarkonia. The observed agreement demonstrates that the leading Fock sector LFWFs projected from BSA are sufficiently accurate to describe the hadron structure. Actually, direct calculation of the normalization indicates that the valance Fock sector is dominant for heavy quarkonia. However, for light mesons like the pion, higher Fock states are more important.
@@ Acknowledgements.
We wish to thank X. Zhao for fruitful discussions. This work is supported in part by the National Natural Science Foundation of China (NSFC) under Grant No. 12375081, and by the Chinese Academy of Sciences under Grant No. YSBR-101.

### v3

# Convergence in charmonium structure: light-front wave functions from basis light-front quantization and Dyson-Schwinger equations

@@ Abstract
The charmonium system, bound states of charm and anti-charm quarks, serves as a critical laboratory for exploring non-perturbative quantum chromodynamics (QCD) [1]. It is relatively simple compared to light hadrons – owing to the heavy quark mass, making it an ideal testbed for theoretical frameworks that bridge perturbative and non-perturbative physics.
@@ I Light-front wave functions
where, $p$ is the 4-momentum of the particle; $j$ and $\sigma$ are the particle’s total angular momentum and magnetic projection, respectively. The coefficients $\psi_{s\bar{s}/h}^{\sigma}(x,\vec{k}_{\perp})$ are the valence sector LFWFs, with $s(\bar{s})$ representing the spin of the quark (antiquark). $x=p^{+}_{1}/p^{+}$ is the longitudinal momentum fraction, $\vec{k}_{\perp}=\vec{p}_{1\perp}-x\vec{p}_{\perp}$ is the relative transverse momentum and $p_{1}^{2}=p_{2}^{2}=m_{q}^{2}$. The ellipsis represents the non-valence Fock contributions. The quark and antiquark creation operators satisfy the anticommutation relations:
@@
$\psi_{\uparrow\downarrow-\downarrow\uparrow/P}(x,\vec{k}_{\perp})=\frac{1}{2p^{+}}\int{\mathrm{d}}^{3}z\,e^{{\frac{i}{2}}xp\cdot z-i\vec{k}_{\perp}\cdot\vec{z}_{\perp}}\\ \times\langle 0|\bar{\psi}(0)\gamma^{+}\gamma_{5}\psi(z)|P(p)\rangle_{z^{+}=0},$ (4)
@@
The second term appears because the LFWFs in general depend on the orientation of light front $\omega$, which is a null vector ($\omega^{2}=0$, and $\omega\cdot\gamma=\gamma^{+}$). In some phenomenological applications, the second term is dropped [64, 65, 66, 67, 68, 69, 70]. However, we have shown that this term plays an important role in chiral symmetry breaking [71]. In both BLFQ and DSE, the maximal kinematical symmetries of the LFWFs are retained and both approaches produce two independent structures. Figure 1 compares the valence LFWFs of $\eta_{c}$ as obtained from BLFQ and from DSE. The LFWFs of BLFQ and DSE are accessible from data repository [72] and [73], respectively. For both spin-singlet $\psi_{\uparrow\downarrow-\downarrow\uparrow}$ and spin-triplet $\psi_{\uparrow\uparrow}/\psi_{\downarrow\downarrow}$ components, the LFWFs from BLFQ are broader than DSE in the transverse momentum ($k_{\perp}$) direction and longitudinal ($x$) direction. However, the normalizations of each of the spin configuration are qualitatively similar in these two frameworks. The normalization for each spin configuration is defined as
@@
A comment is in order here. The valence Fock sector LFWFs projected from the DSE are not automatically normalized to 1 due to contributions from high Fock states in the BSA. However, for heavy quarkonia studied in this paper, the valence Fock sector contributes more than 90% of the total normalization [54]. To be consistent with the BLFQ approach, we have normalized the DSE LFWFs to 1. All subsequent calculations will use these normalized LFWFs.
In practical calculations, the LFWFs within the BLFQ are expanded in a basis of 2D harmonic oscillator functions in the transverse directions and Jacobi functions in the longitudinal momentum fraction $x$ [21]. To render the basis finite for numerical implementation, the quantum numbers are truncated according to $2n+|m|+1\leq N_{\text{max}}$ and $0\leq l\leq L_{\text{max}}$, where $n,m$ denote the radial and angular quantum numbers of the 2D harmonic oscillator modes and $l$ labels the order of the Jacobi functions. This $N_{\text{max}}$-truncation naturally introduces both UV and infrared IR cutoffs: $\Lambda_{\text{UV}}=b\sqrt{N_{\text{max}}}$, $\lambda_{\text{IR}}=b/\sqrt{N_{\text{max}}}$, where $b$ is the oscillator basis scale. Meanwhile, $L_{\text{max}}$ controls the resolution in the longitudinal momentum fraction $\Delta x\approx L_{\text{max}}^{-1}$. In contrast, the DSE approach adopts an explicit UV cutoff $\Lambda$ on the Euclidean momentum $k_{E}$ of the Green’s functions. In practice, $\Lambda$ is chosen sufficiently large to ensure convergence while minimizing regulator artifacts [54].
@@ II Form factors and hadronic densities
$\langle O\rangle=\sum_{n}\frac{1}{S_{n}}\sum_{s_{i}}\prod_{i}^{n}\int\frac{{\mathrm{d}}x_{i}}{4\pi x_{i}}\int{\mathrm{d}}^{2}r_{i\perp}4\pi\delta(\sum_{j}x_{j}-1)\\ \times\delta^{2}(\sum_{j}x_{j}\vec{r}_{j\perp})\big|\widetilde{\psi}_{n}(\{x_{i},\vec{r}_{i\perp},s_{i}\})\big|^{2}O(\{x_{i},\vec{r}_{i\perp},s_{i}\}).$ (10)
Here, $S_{n}$ is the symmetry factor. $\psi_{n}(\{x_{i},\vec{r}_{i\perp},s_{i}\})$ is the transverse coordinate space wave function defined as,
$\widetilde{\psi}_{n}(\{x_{i},\vec{r}_{i\perp},s_{i}\})=\prod_{i=1}^{n}\int\frac{{\mathrm{d}}^{2}k_{i\perp}}{(2\pi)^{2}}(2\pi)^{2}\delta^{2}(\sum_{i}k_{i\perp})\\ \times e^{-i\sum_{i}\vec{k}_{i\perp}\cdot\vec{r}_{i\perp}}\psi_{n}(\{x_{i},\vec{k}_{i\perp},s_{i}\}).$ (11)
For charmonium, the physical charge form factor vanishes due to the charge conjugation symmetry. Nevertheless, it is custom to define a fictitious charge form factor, where the photon couples to the quark and antiquark differently, similar to the charged pion $\pi^{\pm}$ and $B_{c}$, while the quark masses remain unchanged $m_{q}=m_{\bar{q}}=m_{c}$ [80]. Figure 2 compares the charge form factor computed using LFWFs from BLFQ and from DSE. Following our previous analyses for dilepton, diphoton and radiative transitions, hereafter we adopt $N_{\mathrm{max}}=8$ LFWFs for central values of BLFQ results, which corresponds to a UV scale $\Lambda_{\text{UV}}=2.8\,\text{GeV}\approx M_{c\bar{c}}$, and quote the difference between $N_{\mathrm{max}}=8$ and $N_{\mathrm{max}}=16$ as the basis sensitivity [21, 27, 28], where $N_{\mathrm{max}}$ denotes the basis truncation parameter. This choice is motivated by the competition between the needs for both a better basis resolution and a lower UV scale to suppress the radiative corrections. The primary source of uncertainty in the DSE results stems from the RL approximation, which is difficult to estimate. Consequently, only the central values are presented here and in the subsequent calculations. As shown in Fig. 2, the BLFQ and DSE results are in good agreement with each other including at high $Q^{2}$ ($Q^{2}\gtrsim 10\,\mathrm{GeV^{2}}$).
@@
$\langle\psi(p^{\prime})|T^{\mu\nu}(0)|\psi(p)\rangle=2P^{\mu}P^{\nu}A(Q^{2})\\ +\frac{1}{2}(q^{\mu}q^{\nu}-g^{\mu\nu}q^{2})D(Q^{2}).$ (12)
In light-front dynamics, $A$ and $D$ are related to the Fourier transform of the transverse flow density $\mathcal{A}(r_{\perp})$, and transverse shear density $\mathcal{T}^{12}(r_{\perp})=\mathcal{S}(r_{\perp})/(2P^{+})$
@@
Figure 3 compares the gravitational form factors $A(Q^{2})$ and $D(Q^{2})$ computed using LFWFs from BLFQ and from DSE. As one can see, the BLFQ and DSE results are in good agreement. The $D$-term, defined as $D\equiv D(0)$, is known to be sensitive to the interaction. In Ref. [85], the authors argue that the $D$-term of the pseudoscalar meson should lie within (1,1/3)(-1,-1/3), which are the values in the chiral limit and the infinite mass limit respectively. Using DSE with a contact interaction, they obtain D=0.58D=-0.58 for $\eta_{c}$. In contrast, the values extracted from BLFQ and from DSE LFWFs in this work are $D_{\textsc{blfq}}=-4.6$ and $D_{\textsc{dse}}=-4.1$.
@@ III Light-cone distribution amplitudes and radiative transitions
$\langle 0|\bar{\psi}(-{\frac{1}{2}}z)\gamma^{+}\gamma_{5}\psi(+{\frac{1}{2}}z)|P(p)\rangle_{\mu}\\ =if_{P}p^{+}\int_{0}^{1}{\mathrm{d}}xe^{{\frac{i}{2}}xp^{+}z^{-}}\phi_{P}(x,\mu).$ (17)
@@
Figure 4 compares the leading-twist LCDA of $\eta_{c}$ as obtained from BLFQ and from DSE and evolved to $\mu=2.8\,\mathrm{GeV}$. The DSE result appears narrower than the BLFQ result. However, both are broad enough to distinguish from the nonrelativistic limit. The decay constant can be derived after integrating over $x$ in Eq. (18),
@@
Figure 4: (Colors online) Comparison of the leading-twist LCDA of $\eta_{c}$ as obtained from BLFQ ($N_{\text{max}}=8$, solid lines) and from DSE (dashed lines). The uncertainty bands are computed from the difference between $N_{\text{max}}=8$ and $N_{\text{max}}=16$ results in BLFQ. The dot-dashed lines correspond to the perturbative QCD (pQCD) asymptotic form 6x(1x)6x(1-x) [7]. The DSE results are evolved from $\mu_{0}=2.6\,\mathrm{GeV}$ to $\mu=2.8\,\mathrm{GeV}$ to match the BLFQ energy scale.
The LCDA is an essential tool for describing exclusive processes in high-energy scattering $Q^{2}\gg\Lambda_{\textsc{qcd}}^{2}$. One of the observables that can be described by LCDA is the two-photon transition form factor (TFF), as shown by Lepage & Brodsky [7]. On the other hand, the evaluation of the TFF can be extended to low $Q^{2}$ via LFWFs [7, 87, 27]. Figure 5 compares the evaluation of the TFF $F_{\eta_{c}\gamma}(Q^{2})$ for the process $\gamma^{*}\gamma\to\eta_{c}$ using LFWFs obtained from BLFQ and from DSE. The theoretical results, without tuning of parameters, are in good agreement with BaBar measurements.
@@ IV Summary and conclusions
This work also validates the reliability of the BSA projection method in heavy quarkonia. The observed agreement demonstrates that the leading Fock sector LFWFs projected from BSA are sufficiently accurate to describe the hadron structure. Actually, direct calculation of the normalization indicates that the valence Fock sector is dominant for heavy quarkonia. However, for light mesons like the pion, higher Fock states are more important.
@@ Acknowledgements.
We wish to thank X. Zhao for fruitful discussions. This work is supported in part by the National Natural Science Foundation of China (NSFC) under Grant No. 12375081, and by the Chinese Academy of Sciences under Grant No. YSBR-101. X.-h. Cao is supported by the NSFC under Grant No. 125B2111.
Data availability
The data that support the findings of this article are openly available [72, 73].